OpenStudy (anonymous):

What is the solution to 4x²+24x+36=0

OpenStudy (anonymous):

This should be pretty straight forward to factor. What's causing the problem?

OpenStudy (anonymous):

4(x^2+6x+9)=0

OpenStudy (anonymous):

then just factor it

OpenStudy (anonymous):

i do not understand how to factor the x and the x²

OpenStudy (anonymous):

To factor you are going to undo the foil process. \[x^2 + (a+b)x + ab = (x+a)(x+b)\]

OpenStudy (anonymous):

So you are looking for two numbers (a and b) that will add together to give you 6, and multiply to make 9.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Did you figure out which numbers you have for a and b?

OpenStudy (anonymous):

ummm.... no, this is hard